Chief SOHCAHTOA decided his cone-shaped teepee needed to be 7 feet tall. He wanted his teepee to have a diameter of 12 feet. How long does he need to make the sides of his teepee.
Answer by stanbon(75887)

Chief SOHCAHTOA decided his cone-shaped teepee needed to be 7 feet tall. He wanted his teepee to have a diameter of 12 feet. How long does he need to make the sides of his teepee.
-------------
Draw a vertical slice of the teepee as a right triangle with base of 6 and
height of 12. Find the hypotenuse using Pythagoras.
hy^2 = 6^2 + 12^2
h^2 = 36 + 144
h^2 = 180
h = 6sqrt(5) ft
-------------------
Cheers,
Stan H.